OpenStudy (anonymous):

A decorator plans to place a rug in a room 9m by 12m in such a way that a uniform strip of flooring around the rug will remain uncovered. If the rug is to cover half the floor space, what should the dimensions of the rug be?

OpenStudy (anonymous):

6m by 9m

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

If the strip is the same width all the way around, the dimensions of the rug will be (9-2x) * (12-2x).

OpenStudy (anonymous):

because that looked to fit the question.... you can also solve it by solving the following: (9-2x)*(12-2x)=9*12*0.5 Which will give you the length of the strip for x

OpenStudy (anonymous):

You want that to be half the area of the 9*12 room, so (9-2x)*(12-2x) = (1/2)(9)(12)

OpenStudy (anonymous):

X is the width of the rugless strip around the room's perimeter